OpenStudy (anonymous):

A swimming pool can be filled in 15 hours if water enters through a pipe alone or in 23 hours if water enters through a hose alone. if water is entering through both the pipe and the hose how long will it take to fill the pool 2/5 full? In simplified form

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

no that's why I cannot figure it out ;(

ganeshie8 (ganeshie8):

Lets say the pool has a volume of "x" liters, and it takes "t" hours to fill the poll if both pipe and hose are open.

ganeshie8 (ganeshie8):

\[\dfrac{x}{15} + \dfrac{x}{23} = \dfrac{x}{t}\] \[\implies t = \dfrac{15\times 23}{15+23} = \dfrac{15\times 23}{38}\]

ganeshie8 (ganeshie8):

^^thats the time it takes to fill the pool `completely`

OpenStudy (anonymous):

I am still confused

ganeshie8 (ganeshie8):

since you want how much time it takes for filling 2/5th of the pool, just take 2/5 of above : Time to fill 2/5th of pool = \(\dfrac{2}{5} \left(\dfrac{15 \times 23}{38}\right) = \dfrac{69}{19} \) hours